OpenStudy (anonymous):

Could someone help me out with this one..? On the graph of the equation 2x − y = 10, what is the value of the y-intercept? -10 -5 5 10

OpenStudy (anonymous):

2x-y=10 y=2x-10 b=-10

OpenStudy (luigi0210):

y-intercept is when x=0 x-intercept is when y=0

OpenStudy (anonymous):

y=mx+b m=slope b=y-intercept

OpenStudy (anonymous):

thanks guys :))

OpenStudy (anonymous):

okay, how do i start this..?

OpenStudy (anonymous):

put the equation into y=mx+b form

OpenStudy (anonymous):

oh, i see

OpenStudy (anonymous):

-y = -2x -10

OpenStudy (anonymous):

and the y intercept would be -10

OpenStudy (anonymous):

yes